# Fair Classifiers Without Fair Training: An Influence-Guided Data Sampling Approach

### v1

@@ Appendix B Omitted Proofs
Following a similar idea, for the second term, we also have
@@
$\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})$ $\leq\underbrace{(\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}}))}_{\text{train-test model gap}}+\underbrace{2L\lVert\mathbf{w}^{k}-\mathbf{w}^{P}\rVert^{2}}_{\text{group gap}}+(\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}}))$
Lastly, integrating with Lemma B.2, B.3 and B.4, we can finish the proof. ∎
Lemma B.2.
@@
[…]
@@
where we define $\text{dist}(\mathcal{P},\mathcal{Q})=\sum_{i=1}^{I}|p^{(\mathcal{P})}(\pi=i)-p^{(\mathcal{Q})}(\pi=i)|$ and $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{P},z)]\leq G_{P},\forall i\in I$ because of Assumption 4.2. For the second term, with probability at least $1-\delta$, we have $|\mathcal{R}_{\mathcal{P}}(\mathbf{w}^{P})-\mathcal{R}_{P}(\mathbf{w}^{P})|\leq\sqrt{\frac{\log(4/\delta)}{2N_{P}}}$. Note that the third term $\mathcal{R}_{P}(\mathbf{w}^{P})-\mathcal{R}_{Q}^{*}(\mathbf{w^{\mathcal{Q}}})$ can be regarded as a constant $\varpi$.because $\mathcal{R}_{P}(\mathbf{w}^{P})$ is the empirical risk and $\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}})$ is the ideal minimal empirical risk of model $\mathbf{w}^{\mathcal{Q}}$ trained on distribution $\mathcal{Q}$.
@@
$\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}})\leq G_{P}\cdot\text{dist}(\mathcal{P},\mathcal{Q})+\sqrt{\frac{\log(4/\delta)}{2N_{P}}}+\varpi.$
Lemma B.3.
(Group gap) Suppose Assumptions 4.1 and 4.2 hold for empirical risk $\mathcal{R}_{P}(\cdot)$, then we have
$\lVert\mathbf{w}^{k}-\mathbf{w}^{P}\rVert^{2}\leq 2LG^{2}\sum_{t=0}^{T}(\eta_{t}^{2}(1+2\eta_{t}^{2}L^{2}))^{t}\bigg{(}\sum_{i=1}^{I}\bigg{|}p^{(k)}(\pi=i)-p^{(P)}(\pi=i)\bigg{|}\bigg{)}^{2}.$
where $\eta_{t}$ is epoch $t$’s learning rate and T is the number of epochs.
@@
$\begin{split}&\lVert\mathbf{w}^{k}_{T}-\mathbf{w}^{P}_{T}\rVert^{2}\\ &\leq(1+2\eta_{t}^{2}L^{2})\lVert\mathbf{w}^{k}_{T-1}-\mathbf{w}^{P}_{T-1}\rVert^{2}+2L\eta_{t}^{2}G^{2}\bigg{(}\sum_{i=1}^{I}|p^{(k)}(\pi=i)-p^{(P)}(\pi=i)|\bigg{)}^{2}\\ &\leq\prod_{t=0}^{T}(1+2\eta_{t}^{2}L^{2})^{t}\lVert\mathbf{w}^{k}_{0}-\mathbf{w}^{P}_{0}\rVert^{2}+2LG^{2}\sum_{t=0}^{T}(\eta_{t}^{2}(1+2\eta_{t}^{2}L^{2}))^{t}\bigg{(}\sum_{i=1}^{I}|p^{(k)}(\pi=i)-p^{(P)}(\pi=i)|\bigg{)}^{2}\\ &\leq 2LG^{2}\sum_{t=0}^{T}(\eta_{t}^{2}(1+2\eta_{t}^{2}L^{2}))^{t}\bigg{(}\sum_{i=1}^{I}|p^{(k)}(\pi=i)-p^{(P)}(\pi=i)|\bigg{)}^{2}.\end{split}$
where the last inequality holds because the initial models are the same, i.e., $\mathbf{w}_{0}=\mathbf{w}^{k}_{0}=\mathbf{w}^{P}_{0},\forall k$. ∎
Lemma B.4.
@@
[…]
@@
Building upon the proof idea presented in Lemma B.2, for completeness, we provide a full proof here. Firstly, we have,
@@
$\begin{split}\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{P}_{k}}(\mathbf{w}^{k})&=\sum_{i=1}^{I}p^{(\mathcal{Q}_{k})}(\pi=i)\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]-\sum_{i=1}^{I}p^{(\mathcal{P}_{k})}(\pi=i)\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]\\ &\leq\sum_{i=1}^{I}|p^{(\mathcal{P}_{k})}(\pi=i)-p^{(\mathcal{Q}_{k})}(\pi=i)|\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]\\ &\leq G_{k}\cdot\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k}).\end{split}$
where $\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k}):=\sum_{i=1}^{I}|p^{(\mathcal{P}_{k})}(\pi=i)-p^{(\mathcal{Q}_{k})}(\pi=i)|$ and $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]\leq G_{k},\forall i\in I$ due to Assumption 4.2. Recall that the constant $G_{k}$ clarifies the bound of loss on the corresponding model $\mathbf{w}^{k}$. For the second term, with probability at least $1-\delta$, we have $|\mathcal{R}_{\mathcal{P}_{k}}(\mathbf{w}^{k})-\mathcal{R}_{P_{k}}(\mathbf{w}^{k})|\leq\sqrt{\frac{\log(4/\delta)}{2N_{P_{k}}}}$. For the third term, we define $\varpi_{k}:=\mathcal{R}_{P_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}})$, which can be regarded as a constant. This is because $\mathcal{R}_{P_{k}}(\mathbf{w}^{k})$ represents empirical risk and $\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w^{\mathcal{Q}_{k}}})$ is the ideal minimal empirical risk of model $\mathbf{w}^{\mathcal{Q}_{k}}$ trained on sub-distribution $\mathcal{Q}_{k}$.
@@
$\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}})\leq G_{k}\cdot\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k})+\sqrt{\frac{\log(4/\delta)}{2N_{P_{k}}}}+\varpi_{k}.$
B.3 Proof of Lemma 5.1
Lemma 5.1. The influence of predictions on the validation dataset $Q_{v}$ can be denoted by
$\textsf{Infl}_{\textsf{acc}}(z^{\prime}):=\sum_{n\in[N_{v}]}\textsf{Infl}_{\textsf{acc}}(z^{\prime},z_{n}^{\circ})\approx-\eta\left\langle\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z^{\prime}),\sum_{n=1}^{N_{v}}\partial_{\mathbf{w}_{t}}\ell(\mathbf{w},z_{n}^{\circ})\right\rangle$
Proof.
Taking the first-order Taylor expansion, we will have
$f(\mathbf{w}_{t+1},x_{n}^{\circ})-f(\mathbf{w}_{t},x_{n}^{\circ})\approx$ $\left.\left\langle\frac{\partial f(\mathbf{w},x_{n}^{\circ})}{\partial\mathbf{w}},\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\rangle\right|_{\mathbf{w}=\mathbf{w}_{t}}=$ $-\eta\left.\left\langle\frac{\partial f(\mathbf{w},x_{n}^{\circ})}{\partial\mathbf{w}},\frac{\partial\ell(\mathbf{w},z^{\prime})}{\partial\mathbf{w}}\right\rangle\right|_{\mathbf{w}=\mathbf{w}_{t}}.$
where the last equality holds due to Eq. (5). Therefore,
$\textsf{Infl}_{\textsf{acc}}(z^{\prime}):=\sum_{n\in[N_{v}]}\textsf{Infl}_{\textsf{acc}}(z^{\prime},z_{n}^{\circ})\approx-\eta\left\langle\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z^{\prime}),\sum_{n=1}^{N_{v}}\partial_{\mathbf{w}_{t}}\ell(\mathbf{w},z_{n}^{\circ})\right\rangle$
B.4 Proof of Lemma 5.2
Lemma 5.2 The influence of fairness on the validation dataset $Q_{v}$ can be denoted by
$\textsf{Infl}_{\textsf{fair}}(z^{\prime}):=\sum_{n\in[N_{v}]}\textsf{Infl}_{\textsf{fair}}(z^{\prime},z_{n}^{\circ})\approx-\eta\sum_{n\in[N_{v}]}\left\langle\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z^{\prime}),\partial_{\mathbf{w}_{t}}\phi(\mathbf{w}_{t},z_{n}^{\circ})\right\rangle$
Proof.
By first-order approximation, we have
Recall by first-order approximation, we have
Therefore,
$\textsf{Infl}_{\textsf{fair}}(z^{\prime})=\sum_{n\in[N_{v}]}\textsf{Infl}_{\textsf{fair}}(z^{\prime},z_{n}^{\circ})\approx-\left.\eta\sum_{n\in[N_{v}]}\left\langle\frac{\partial\ell(\mathbf{w},z^{\prime})}{\partial\mathbf{w}},\frac{\partial\phi(\mathbf{w}_{t},z_{n}^{\circ})}{\partial\mathbf{w}}\right\rangle\right|_{\mathbf{w}=\mathbf{w}_{t}}.$
Appendix C Exploration of the Labeling Strategies
  1. Strategy I
  2. Strategy II
Remark C.1.
Proof.
Based on the definition of the influence of prediction component, as delineated in Eq. (6), it becomes evident that the most uncertain points are obtained when the proxy labels closely align with the true labels. Consequently, the model predictions used in Strategy II also approximate the true labels in order to minimize the cross-entropy loss. Thus, in a certain sense, Strategy I and Strategy II can be considered equivalent. ∎
Appendix D Additional Experimental Results
D.1 Datasets and Parameter Settings
We empirically evaluate FIS on the CelebA dataset, an image dataset commonly used in the fairness literature [Liu et al., 2015]. We also evaluate FIS on two tabular datasets: UCI Adult [Asuncion and Newman, 2007] and Compas Dataset [Angwin et al., 2016].
D.1.1 CelebA Dataset
@@
In all our experiments using CelebA dataset, we train a vision transformer with patch size $(8,8)$ using SGD optimizer and a batch size of 128. The epochs are split into two phases: warm-up epochs (5 epochs) and training epochs (10 epochs). The default label budget per round, which represents the number of solicited data samples, is set to 256. Additionally, the default values for learning rate, momentum, and weight decay are 0.01, 0.9, and 0.0005, respectively. We initially allocate 2% of the data for training purposes and the remaining 98% for sampling. Then, we randomly select 10% of the test data for validation. For JTT, we explore 10% data for training purposes with weights $\lambda\in\{20,100\}$ for retraining misclassified examples.
D.1.2 UCI Adult Dataset
@@
In the experiments using Adult dataset, we train a two-layer ReLU network with a hidden size of 64. The epochs are split into two phases: warm-up epochs (20 epochs) and training epochs (10 epochs). The default label budget per round, which represents the number of solicited data samples, is set to 1024. Additionally, the default values for learning rate, momentum, and weight decay are 0.00001, 0.9, and 0.0005, respectively. We resample the datasets to balance the class and group membership [Chawla et al., 2002]. The dataset is randomly split into a train and a test set in a ratio of 80 to 20. Then, we randomly re-select 20% of the training set for initial training and the remaining 80% for sampling. Also, 20% examples of the test set are selected to form a validation set. We utilize the whole model to compute the prediction influence and fairness for examples. Then, we randomly select 10% of the test data for validation. For JTT, we explore 30% data for training purposes with weights $\lambda\in\{20,100\}$ for retraining misclassified examples.
D.1.3 Compas Dataset

### v3

@@ Appendix D Omitted proofs
where $(\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}}))$ denotes the ideal minimal empirical risk of model $\mathbf{w}^{\mathcal{Q}}$ trained on distribution $\mathcal{Q}$. Following a similar idea, for the second term, we also have
@@
$\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})$ $\leq\underbrace{(\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}}))}_{\text{train-test model gap}}+\underbrace{2L\lVert\mathbf{w}^{k}-\mathbf{w}^{P}\rVert^{2}}_{\text{group model gap}}+(\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}}))$
Lastly, integrating with Lemmas D.1, D.2 and D.3, we can finish the proof. ∎
Lemma D.1.
@@
$\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}})\leq G_{P}\cdot\text{dist}(\mathcal{P},\mathcal{Q})+\sqrt{\frac{\log(4/\delta)}{2|P|}}+\varpi.$
where $\text{dist}(\mathcal{P},\mathcal{Q})=\sum_{i=1}^{I}|p^{(\mathcal{P})}(\pi=i)-p^{(\mathcal{Q})}(\pi=i)|$ and $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{P},z)]\leq G_{P},\forall i\in I$, and a constant $\varpi:=\mathcal{R}_{P}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}})$.
@@
where we define $\text{dist}(\mathcal{P},\mathcal{Q})=\sum_{i=1}^{I}|p^{(\mathcal{P})}(\pi=i)-p^{(\mathcal{Q})}(\pi=i)|$ and $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{P},z)]\leq G_{P},\forall i\in I$ because of Assumption 5.2. For the second term, with probability at least $1-\delta$, we have $|\mathcal{R}_{\mathcal{P}}(\mathbf{w}^{P})-\mathcal{R}_{P}(\mathbf{w}^{P})|\leq\sqrt{\frac{\log(4/\delta)}{2|P|}}$. Note that the third term $\mathcal{R}_{P}(\mathbf{w}^{P})-\mathcal{R}_{Q}^{*}(\mathbf{w^{\mathcal{Q}}})$ can be regarded as a constant $\varpi$.because $\mathcal{R}_{P}(\mathbf{w}^{P})$ is the empirical risk and $\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}})$ is the ideal minimal empirical risk of model $\mathbf{w}^{\mathcal{Q}}$ trained on distribution $\mathcal{Q}$.
@@
$\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}})\leq G_{P}\cdot\text{dist}(\mathcal{P},\mathcal{Q})+\sqrt{\frac{\log(4/\delta)}{2|P|}}+\varpi.$
Lemma D.2.
(Group model gap) Suppose Assumptions 5.1 and 5.2 hold for empirical risk $\mathcal{R}_{P}(\cdot)$. The initial learning rate $\eta_{0}^{2}<\frac{1}{\sqrt{2}TL}$, where $T$ denotes the number of training epochs. Then, we have
$\lVert\mathbf{w}^{k}-\mathbf{w}^{P}\rVert^{2}\leq 2LG^{2}\bigg{(}\sum_{i=1}^{I}\bigg{|}p^{(k)}(\pi=i)-p^{(P)}(\pi=i)\bigg{|}\bigg{)}^{2}.$
@@
$\begin{split}&\lVert\mathbf{w}^{k}_{T}-\mathbf{w}^{P}_{T}\rVert^{2}\\ &\leq(1+2\eta_{T}^{2}L^{2})\lVert\mathbf{w}^{k}_{T-1}-\mathbf{w}^{P}_{T-1}\rVert^{2}+2L\eta_{T}^{2}G^{2}\bigg{(}\sum_{i=1}^{I}|p^{(k)}(\pi=i)-p^{(P)}(\pi=i)|\bigg{)}^{2}\\ &\leq\prod_{t=0}^{T}(1+2\eta_{t}^{2}L^{2})^{t}\lVert\mathbf{w}^{k}_{0}-\mathbf{w}^{P}_{0}\rVert^{2}+2LG^{2}\sum_{t=0}^{T}(\eta_{t}^{2}(1+2\eta_{t}^{2}L^{2}))^{T-t}\bigg{(}\sum_{i=1}^{I}|p^{(k)}(\pi=i)-p^{(P)}(\pi=i)|\bigg{)}^{2}\\ &\leq 2LG^{2}\sum_{t=0}^{T}(\eta_{t}^{2}(1+2\eta_{t}^{2}L^{2}))^{T-t}\bigg{(}\sum_{i=1}^{I}|p^{(k)}(\pi=i)-p^{(P)}(\pi=i)|\bigg{)}^{2}.\end{split}$
where the last inequality holds because the initial models are the same, i.e., $\mathbf{w}_{0}=\mathbf{w}^{k}_{0}=\mathbf{w}^{P}_{0},\forall k$. When the condition $\eta_{t}^{2}<\eta_{0}^{2}<\frac{1}{\sqrt{2}TL}$ satisfies, $2LG^{2}\sum_{t=0}^{T}(\eta_{t}^{2}(1+2\eta_{t}^{2}L^{2}))^{T-t}$ can be simplifited as $2LG^{2}$, which is independent of the learning algorithm. This condition is easy to be satisfied since the learning rate $\eta_{t}$ is a small value ($<0.0001$) and usually set to be decay with the training epoch (i.e., $\eta_{t+1}\leq\eta_{t}$). ∎
Lemma D.3.
@@
$\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}})\leq G_{k}\cdot\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k})+\sqrt{\frac{\log(4/\delta)}{2|P_{k}|}}+\varpi_{k}.$
where $\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k})=\sum_{i=1}^{I}|p^{(\mathcal{P}_{k})}(\pi=i)-p^{(\mathcal{Q}_{k})}(\pi=i)|$ and $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]\leq G_{k},\forall i\in I$, and $\varpi_{k}:=\mathcal{R}_{P_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}})$.
@@
Building upon the proof idea presented in Lemma D.1, for completeness, we provide a full proof here. Firstly, we have,
@@
$\begin{split}&\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{P}_{k}}(\mathbf{w}^{k})\\ &=\sum_{i=1}^{I}p^{(\mathcal{Q}_{k})}(\pi=i)\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]-\sum_{i=1}^{I}p^{(\mathcal{P}_{k})}(\pi=i)\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]\\ &\leq\sum_{i=1}^{I}|p^{(\mathcal{P}_{k})}(\pi=i)-p^{(\mathcal{Q}_{k})}(\pi=i)|\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]\\ &\leq G_{k}\cdot\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k}).\end{split}$
where $\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k}):=\sum_{i=1}^{I}|p^{(\mathcal{P}_{k})}(\pi=i)-p^{(\mathcal{Q}_{k})}(\pi=i)|$ and $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]\leq G_{k},\forall i\in I$ due to Assumption 5.2. Recall that the constant $G_{k}$ clarifies the bound of loss on the corresponding model $\mathbf{w}^{k}$. For the second term, with probability at least $1-\delta$, we have $|\mathcal{R}_{\mathcal{P}_{k}}(\mathbf{w}^{k})-\mathcal{R}_{P_{k}}(\mathbf{w}^{k})|\leq\sqrt{\frac{\log(4/\delta)}{2|P_{k}|}}$. For the third term, we define $\varpi_{k}:=\mathcal{R}_{P_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}})$, which can be regarded as a constant. This is because $\mathcal{R}_{P_{k}}(\mathbf{w}^{k})$ represents empirical risk and $\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w^{\mathcal{Q}_{k}}})$ is the ideal minimal empirical risk of model $\mathbf{w}^{\mathcal{Q}_{k}}$ trained on sub-distribution $\mathcal{Q}_{k}$.
@@
$\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}})\leq G_{k}\cdot\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k})+\sqrt{\frac{\log(4/\delta)}{2|P_{k}|}}+\varpi_{k}.$
Appendix E More experimental results
E.1 Datasets and parameter settings
We empirically evaluate FIS on the CelebA dataset, an image dataset commonly used in the fairness literature [49]. We also evaluate FIS on two tabular datasets: UCI Adult [8] and Compas dataset [6].
E.1.1 CelebA dataset
@@
In all our experiments using the CelebA dataset, we train a vision transformer with patch size $(8,8)$ using SGD optimizer and a batch size of 128. The epochs are split into two phases: warm-up epochs (5 epochs) and training epochs (10 epochs). The default label budget per round, which represents the number of solicited data samples, is set to 256. Additionally, the default values for learning rate, momentum, and weight decay are 0.01, 0.9, and 0.0005, respectively. We initially allocate 2% of the training set for training purposes and the remaining 98% for sampling. Then, we randomly select 10% of the test data for validation. For JTT, we explore 10% data for training purposes with weights $\lambda=20$ for retraining misclassified examples.
E.1.2 UCI Adult dataset
@@
In the experiments using the Adult dataset, we train a two-layer ReLU network with a hidden size of 64. The epochs are split into two phases: warm-up epochs (100 epochs) and training epochs (60 epochs). The default label budget per round, which represents the number of solicited data samples, is set to 1024. Additionally, the default values for learning rate, momentum, and weight decay are 0.00001, 0.9, and 0.0005, respectively. We resample the datasets to balance the class and group membership [17]. The dataset is randomly split into a train and a test set in a ratio of 80 to 20. Then, we randomly re-select 20% of the training set for initial training and the remaining 80% for sampling. Also, 20% examples of the test set are selected to form a validation set. We utilize the whole model to compute the prediction influence and fairness for examples. Then, we randomly select 10% of the test data for validation. For JTT, we explore 30% data for training purposes with weights $\lambda=20$ for retraining misclassified examples.
E.1.3 Compas dataset